OpenStudy (anonymous):

what is x equal to if ln(lnx)=1 ?

OpenStudy (freckles):

if x=y then e^x=e^y

OpenStudy (freckles):

\[\text{ if } \ln(\ln(x))=1 \text{ then } e^{\ln(\ln(x))}=e^1 \]

OpenStudy (freckles):

\[\text{ and } e^{\ln(y)}=y \text{ for } y > 0\]

OpenStudy (anonymous):

thanks! this came up on my test and i think i got it right